OpenStudy (anonymous):

Solve for x: |x-4| < |x+2|

OpenStudy (clongoria2):

What is the equation?

OpenStudy (anonymous):

here ya go

OpenStudy (anonymous):

I know how to solve it through graphing

OpenStudy (anonymous):

What about the algebra-way?

OpenStudy (anonymous):

well x could equal 6 so it becomes 6-4=2<6+2=8

OpenStudy (clongoria2):

@UltimateWitherLord Couldn't it also be 1 since a variable alone is usually considered 1 Given the fact that 1 - 4 = 3- and 1 + 2 = 3 the answer would still apply

OpenStudy (anonymous):

Answer is square everything!